OpenStudy (anonymous):

What is the maximum velocity (in m/s) acquired by an alpha particle when accelerated between two plates with a potential difference of 580.0 volts? Note: An alpha particle is essentially the nucleus of a helium atom - 2 protons and 2 neutrons.

OpenStudy (abb0t):

I think you can use the formula: \(\sf \color{blue}{\frac{kQ}{r^2}}\) where "k" is the constant: \(\sf 9.0 \times 10^9 \frac{N~ m^2}{C^2}\)

OpenStudy (anonymous):

I need a step by step process to help me with this question. Thank you.